# Very Basic Set theory

### v3

@@
(Date: August 18, 2024)
George Boolos describes the naive conception of set: ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads:
@@ 1. What are individuals?
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains:
@@
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304):
@@
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”:
@@
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $t$ contains $s$ or $s$ is a member of $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ ($s\in t$ or $s\notin t$) but never ($s\in t$ and $s\notin t$). If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox.
@@
no answer: If $s\in s$ then $s\notin s$ and vice versa. Thus $s$ is nonexisting. Therefore:
”No set can contain all and only those sets which do not contain themselves”. $\Box$
The membership relation is definite thus for every $s$ ($s\in s$ or $s\notin s$).
@@
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$) therefore:
For every $s$ exists $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
That is logically equivalent to Russell’s paradox:
It is not the case that exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
Thus for every $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$. $\Box$
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220):
@@
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215):
@@
Then ”which can be thought of as one,” a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’ rather than ‘contained in a set’.
Set theory is a language that follows the rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility. Individuals are possible. Then, hopefully, someone reading or listening that language forms a correct mental picture in her/his mind. That is called ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains:
@@ 2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. There is a membership relation between sets.
Very basic set theory is based on
the world of individuals,
the membership relation is definite,
@@
3. Very basic set theory
@@
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (‘not $\Phi$’) and conjunction $\Phi\wedge\Psi$ (‘$\Phi$ and $\Psi$’). Negation has higher precedence than conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
@@
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set $s$ $\Phi$’, and $\exists s(\Phi)$ means ‘exists set $s$ such that $\Phi$’. Then e.g. $\neg\exists s\forall u(\Phi)$ is short for the formula $\neg(\exists s(\forall u(\Phi)))$.
@@
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. If $s$ is free in $\Phi$ and $\neg\exists s(\Phi\wedge\neg\Phi)$ then ‘$\Phi$ is definite’.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$. The precedence of each of it is lower than conjunction.
$\Phi\vee\Psi$ (‘$\Phi$ or $\Psi$ or both’) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ (‘$\Phi$ implies $\Psi$’) is short for the formula ($\neg\Phi\vee\Psi$),
$\Phi\leftrightarrow\Psi$ (‘$\Phi$ if and only if $\Psi$’) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
If $\Phi$ and $\Psi$ are definite then $\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
@@
Then Russell’s paradox is the formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, and is logically equivalent to $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
The naive conception of set is inconsistent: formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is not definite. Therefore we postulate ‘specification’ (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then $(u\in s\wedge\Phi)$ is definite and
    (specification) $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
@@
We prove: If $\exists u(u\in s\wedge u\in u)$ then exists $v$ whose members are all and only
those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to $s$ to find $v$
@@
We prove ‘the set of all sets’ is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties.
Apply ‘specification’ to $s$ to find $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a member of $v$.
@@
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and $v$ is not a member of $s$. Therefore ‘the set of all sets’ is nonexisting. $\Box$
4. More Axioms
Experience shows a set is equal to or a set is not equal to another set.
We define equality for sets.
@@ 4. More Axioms
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists $\{s,t\}$ and there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
@@
  • (regularity) $\forall s\exists v(v\in s\wedge(v\notin v\rightarrow\forall u((u\in s\wedge u\in v)\rightarrow u\in u)))$.
@@
$\forall v(v\in s\wedge v\notin v\rightarrow\exists u(u\in s\wedge u\in v\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous one $t\in s$, $u\in t$, $v\in u$ is nonexisting. Thus $(t\in s\wedge s\in t)$ is nonexisting therefore $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
@@
An ‘infinite’ sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is nonexisting.
Any pair of different consecutive sets $\{w,\bigcup w\}$ implies $\exists v(v\in w\wedge v\notin v)$ otherwise $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\ldots$ then
$\forall w((w\in z\wedge w\notin w)\rightarrow\exists v(v\in w\wedge v\in\bigcup z\wedge v\notin v))$.
Then $\forall v(v\in\bigcup z\rightarrow v\in z)$ thus $z$ contradicts regularity.
Therefore the sequence terminates: $\exists w(\bigcup w=w)$. Set $z$ is called ‘the transitive closure’ of $s$. Thus for every set the transitive closure exists. $\Box$
5. Conclusion
Very basic set theory is a set theory with individuals
the only constituent of an individual is that individual itself,
Russell’s paradox is logically equivalent to a formula with individuals,
the conjunction of mutually exclusive properties does not constitute a set,
the naming of individuals is logically first,
for every set the transitive closure exists.
Compared with ‘mainstream set theory’ a big difference is ‘what is logically first’. The logic of set theory allows for individuals, and Russell’s paradox is logically equivalent to a formula with individuals. Therefore we choose for ‘the naming of individuals is logically first’, not for the postulate of an ‘empty set’.
There are more differences. The naive conception of set is inconsistent thus axiom ‘specification’ is modified. The axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow\forall t(t\in u\rightarrow t\in s))$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every $s$,
Thus Cantor’s Theorem does not hold.
The definition of equality and axiom ‘pairs’ are only meaningful in combination with sets that are logically first. Thus with axiom ‘regularity’. And ‘regularity’ is used with ‘union’ to prove that for every set the transitive closure exists.
In the last section we define ‘number sequences’ satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a ‘first natural number’. Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, a set of all natural numbers with a specified first number is not excluded.
If in the world only one pair of individuals exists then indeed a set of all natural numbers with that pair of individuals as first number can postulated to exist. That is the way of ‘mainstream set theory’: There is no need for individuals, and the ‘empty set’ is postulated as ‘first natural number’.
@@ 6. Natural numbers
If $u$ and $v$ are the only individuals belonging to transitive set $s$ then $\{u,v\}$ is the only possible set of individuals belonging to $s$. Therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals belonging to $s$ or to $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
@@
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $s\notin s\rightarrow\bigcup s\notin\bigcup s$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. Natural numbers are also known as ‘finite ordinals’.
@@
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number (in physical reality called a ‘unit of measurement’) $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number (unit of measurement) $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence or to the $T$ sequence then $x\in u$ implies $u$ is an $x$ coordinate and $y\in u$ implies $u$ is an $y$ coordinate.

### v4

@@
(Date: January 15, 2025)
George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads
@@ 1. What are individuals?
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains
@@
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304)
@@
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”
@@
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is ‘contained’ in $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ either $s\in t$ or $s\notin t$ but not both. If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox
@@
no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
For every $s$ and for every $t$ the membership relation $s\in t$ is definite. Thus
@@
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$). Then for every $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$.
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220)
@@
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215)
@@
Then ”which can be thought of as one,” thus a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility, and $s\in s$ is possible. Then someone reading or listening that language forms a mental picture of a set in her/his mind. The picture in her/his mind agrees or disagrees with the picture in your mind. We call that ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains
@@ 2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. Therefore set theory describes
a world of individuals,
with a definite membership relation between sets,
@@
Set theory follows rules of logic but the membership relation is definite. That means the logic of ‘true’ and ‘false’ should be changed to the logic of ‘definiteness’.
In the next section we prove logical formulas preserve definiteness.
3. Logic and definiteness
@@
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$). Negation takes precedence over conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
If it is not the case that both $\Phi$ and $\neg\Phi$, that is formula $\neg(\Phi\wedge\neg\Phi)$, then ‘$\Phi$ is definite’. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus formula $s\in t$ is definite.
@@
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’. The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’. The parenthesis following $\forall s$ and $\exists s$ are mandatory. However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
Set $s$ in $\forall s$ or in $\exists s$ is not a ‘variable’. $\forall s$ means ‘for every set’, thus also for a set denoted by $s$. And $\exists s$ means ‘exists a set’ denoted by $s$. Set theory is a language.
@@
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise exists $s$ such that $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ and $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $(\Phi\wedge\neg\Phi)$.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula ($(\neg\Phi\vee\Psi)$,
$\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
There are many rewite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite.
The logical ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then
$\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$, $\Phi\leftrightarrow\Psi$, $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite.
Therefore the building up of formulas preserves definiteness.
4. Russell’s paradox
@@
We prove formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, that is Russell’s paradox
Assume exists set $s$ such that for every set $u$ $(u\in s\leftrightarrow u\notin u)$.
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that $(s\in s\leftrightarrow s\notin s)$.
Formula $(s\in s\leftrightarrow s\notin s)$ is short for $(s\in s\vee s\in s)\wedge(s\notin s\vee s\notin s)$ thus
$(s\in s\wedge s\notin s)$.
Formula $s\in s$ is definite thus it is not the case that $(s\in s\wedge s\notin s)$.
Therefore set $s$ does not exists.
Formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is definite therefore $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$. $\Box$
Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin u\wedge u\notin u))$.
Therefore the logic of definiteness proves that
for every set $s$, an individual is a constituent of $s$ or a set that is not an individual
is not a constituent of $s$.
5. Axiom ‘specification’
If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$. Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$. That is axiom specification (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
@@ 5. Axiom ‘specification’
We prove: If an individual is a constituent of $s$ then exists $v$ whose constituents
are all and only those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
@@
We prove: The set of all sets is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
@@
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and not a constituent of $s$. Therefore the set of all sets is nonexisting. $\Box$
6. Axioms for ‘equality’
@@ 7. Axiom ‘pairs’
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
Therefore we postulate ‘regularity’.
8. Axiom ‘regularity’
@@
  • (regularity) $\forall s\exists v(v\in s\wedge\forall u((u\in s\wedge u\in v)\rightarrow u\in u))$.
@@
$\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
$t\in s$, $u\in t$, $v\in u$ is nonexisting.
Thus $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
9. Axiom ‘union’
@@
We prove: Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
$\bigcup s\neq\bigcup(\bigcup s)\rightarrow\exists u\exists t(u\in t\wedge t\in s\wedge u\notin u\wedge t\notin t)$ otherwise $\bigcup s=\bigcup(\bigcup s)$,
$\bigcup(\bigcup s)\neq\bigcup(\bigcup(\bigcup s))\rightarrow\exists v\exists u\exists t(v\in u\wedge u\in t\wedge t\in s\wedge v\notin v\wedge u\notin u\wedge t\notin t)$,
Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
Therefore the sequence terminates thus exists $w$ belonging to the sequence and $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$. Set $z$ is called ‘the transitive closure’ of $s$. For every set the transitive closure exists. $\Box$
The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and axiom regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘natural numbers’.
@@ 10. Natural numbers
In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”. In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
@@
If the only individuals that are constituent of transitive set $s$ are $u$ and $v$ then $\{u,v\}$ is the only possible set of individuals that is a member of $s$ therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals that are members of $s$ or of $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
@@
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $(s\notin s\rightarrow\bigcup s\notin\bigcup s)$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a first natural number. Natural numbers are also known as ‘finite ordinals’.
@@
Then $\alpha$ is a natural number. In physical reality a first number is called ‘the unit of measurement’. The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence then $x\in u$ implies $u$ is an $x$ coordinate, if $v$ belongs to the $T$ sequence then $y\in v$ implies $v$ is an $y$ coordinate. Then $\{u,v\}$ is an ‘ordered pair’ with an intrinsic intuitive meaning. The Kuratowski ordered pair does not have any intrinsic intuitive meaning. It is ”just a convenient way” to define ordered pairs (Mendelson [2015] p.235, Kanamori [2003] §5).
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded. Then $\bigcup\omega=\omega$. And if $\bigcup\beta=\beta$ we can define ‘transfinite ordinal numbers with first number $\beta$’ analogous to natural numbers (finite ordinal numbers) with first number $\alpha$.
In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
Thus the logic in Very basic set theory is adequate.
Conclusion
Very basic set theory is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first. But individuals are ‘not needed for mathematical purposes’ (Oliver and Smliley [2006] §5.1). Thus Russell’s paradox is not the result of the logic of definiteness otherwise individuals exist, and the naming of a special set is logically first (an ‘empty set’, or an ‘infinite set’). Individuals are banished from ‘mainstream set theory’.
There are more differences with ‘mainstream set theory’. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every set $s$,
Thus Cantor’s Theorem does not hold.